Joe's father is 4 times as old as joe.6 years from now, Joe's father will be 3 times as old as Joe. How old is each now?
Answer by Theo(13342)

x = joes's age now.
y = joes's father's age now.
today, joes's father's age is 4 times joe's age.
you get:
y = 4x
in 6 years, joe's father's age will be 3 times joe's age.
you get:
y + 6 = 3 * (x + 6)
simplify to get:
y + 6 = 3 * x + 18
subtract 6 from both sides of he equation to get:
y = 3x + 12
you have two equations that need to be solved simultaneously.
they are:
y = 4x
y = 3x + 12
since y = 4x in the first equation, replace y with 4x in the second equation to get:
4x = 3x + 12
subtract 3x from both sides of the equation to get:
x = 12
that would be joe's age now.
that makes joe's father's age 4 * 12 = 48 now.
in 6 years, joe will be 12 + 6 = 18
in 6 years, joe's father will be 48 + 6 = 54
54 is equal to 3 * 18.
requirement of the problem are satisfied.
the solution is:
joe is 12 years old and joe's father is 48 years old now.